The electrical breakdown in air occurs at electric field strength of 3.5MV/m
.
Calculate the corresponding intensity of a plane wave.
1
Expert's answer
2019-08-14T14:20:28-0400

We can find the corresponding intensity of a plane wave from the formula:


"I = \dfrac{1}{2}c\epsilon_0E^2,"

here, "c = 3 \cdot 10^8 m/s" is the speed of light, "\epsilon_0 = 8.85 \cdot 10^{-12} F/m" is the permittivity of free space and "E = 3.5 \cdot 10^6 V/m" is the electric field strength.

Finally, we can calculate the corresponding intensity of a plane wave:

"I=\dfrac{1}{2} \cdot 3 \cdot 10^8 \dfrac{m}{s} \cdot 8.85 \cdot 10^{-12} \dfrac{F}{m} \cdot (3.5 \cdot 10^6 \dfrac{V}{m})^2 = 1.63 \cdot 10^{10} \dfrac{W}{m^2}."

Answer:

"I = 1.63 \cdot 10^{10} \dfrac{W}{m^2}."